Imperial Jewelers manufactures and sells a gold bracelet for $403.00. The companys accounting system says that the unit prod• отоптаст“ сорау ассРе сте Реста отас Complete this question by entering your answers in the tabs below. Required 1 Requir
Answer #1

Special order revenue = 16*363 = 5808

Relevant cost of making = (142+6+85+7)*16+460 = 4300

Financial advantage = 1508

2. Yes